A man uses a pulley system to raise a 150lb load to a height of 10ft. if he exerts a force on the rope
50lb through 35ft to accomplish the work;
a. How much work does he do?
b. What is the efficiency of the machine?
1
Expert's answer
2018-07-02T09:22:08-0400
W=Fg=(150)(10)=1500 ft lb.
η=1500/(50(35))=0.857 or 85.7%.